5 numbers chosen randomly without replacement. "B" represents number of even numbers, this random variable has this probability:

x 0 1 2 3 4 5
p(B=x) 0.02693 0.15989 .33858 .31977 .13464 .02020

number of odd #s chosen would then be 5-x, if x is even #s chosen. "C" represents difference b/w # of even and # of odd chosen, --> C= 2B-5

a. probability that exactly 1 even # chosen?

b. probability at most 1 even # chosen?

c. prob. that more even than odd #s chosen?

d. show that mean of B is =2.4359 .
Answer #1

Please don't hesitate to give a "thumbs up" for the answer in case the answer has helped you

a. P(X=1) = .15989

b. P(X<=1) = P(X=0)+P(X=1) = .02693+.15989 = 0.18682

c. P(more even than odd) = P(X=3) + P(X=4) + P(X=5) = .31977+.13464+.02020 = 0.47461

d. The mean is x weighted by p(B=x) = 0*0.02693 + 1*0.15989 +2*.33858+3*.31977+4*.13464+5*.02020 = 2.4359 <Hence Proved>